OpenStudy (anonymous):

Solve 5x^2 = –45x and show how you got that answer?

myininaya (myininaya):

add 45x on both sides then factor

OpenStudy (anonymous):

5x^2 = –45x 5x^2+45x=0 x(5x+45)=0 x=0 or 5x+45=0 5x=-45 x=-9